OpenStudy (anonymous):

1/4r+1/16+1/2r=1/2+r solve the equation

OpenStudy (anonymous):

Easiest route is to find a common denominator. Is the r in the denominator or numerator?

OpenStudy (anonymous):

r is beside it like this 

OpenStudy (anonymous):

\[\frac{4r}{16}+\frac{1}{16}+\frac{8r}{16}=\frac{8}{16}+r\]\[\frac{12r}{16}-\frac{7}{16}=r\]\[12r-7=16r\]\[r=-1.75\]Sorry for the long wait

OpenStudy (anonymous):

thanks